# Towards Understanding the Mechanisms of Classifier-Free Guidance

### v1

@@ F.2 CFG in the Linear Regime
We provide additional experimental results for section 4.1 in Figures 28, 30, 29 and 31. Because the precise transition time from the linear to the nonlinear regime—as well as the influence of each CFG component—varies across classes, we empirically choose the interval for applying guidance and calculate the FDDINOv2 score with 50,000 generated images for each class separately. We summarize our observations as follows (see also section 4.1):
@@
  • Role of CPC Guidance. CPC guidance generally improves generation quality, though its benefits can sometimes be less pronounced. For instance, in the tree frog and castle classes (Figure 30), the CPC term does not enhance FDDINOv2 as much as the mean-shift term. Nevertheless, CPC guidance operates effectively over a wider range of guidance strengths $\gamma$ and noise intervals. For the green mamba and basketball classes, we show results within the prescribed noise interval as solid curves, and extend beyond this interval as dashed curves. While mean-shift becomes highly detrimental once outside the linear regime, CPC guidance remains beneficial.
@@ F.3 Mean-Shifted Noise Initialization
We perform the above experiments on 5 classes, where $\sigma(T)$ is set to 31.9. The results are shown in Figures 32, 33, 34, 35 and 36. Note that the sample quality improves with a properly chosen $\gamma$.
@@ F.4 CFG in the Nonlinear Regime
We provide additional experimental results for section 4.2 in Figure 37. We argue that effective guidance in this regime should satisfy two key properties:
  • Capture local structure of a specific sample. As shown in Figure 26, when $\sigma(t)$ is small, the model diverges considerably from its linear approximation, and linear CFG deviates from the actual nonlinear CFG. In this regime, CFG does not alter the overall image structure but instead refines existing features to produce crisper images. Consequently, effective guidance must adapt to each specific sample. We propose that such guidance can be derived from the network Jacobians $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t),\bm{c})$ evaluated at $\bm{x}_{t}$. Prior work [32] shows that the singular vectors of these Jacobians, which are equivalent to the posterior covariances, adapt to the input $\bm{x}_{t}$.
  • Capture class-specific patterns. As in the linear case, the guidance must also capture class-specific patterns. This can be achieved by contrasting the conditional Jacobian $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t),\bm{c})$ with the unconditional Jacobian $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t))$. Figure 37 shows that guidance built using CPCs—i.e., the difference between these two Jacobians—yields effects similar to actual CFG. In contrast, guidance derived solely from the conditional Jacobian does not improve image quality.
@@
Appendix G CFG in Gaussian Mixture Model
@@
Appendix H Computing Resources

### v3

@@ F.2 CFG in the Linear Regime
We provide additional experimental results for section˜4.1 in Figures˜28, 30, 29 and 31. Because the precise transition time from the linear to the nonlinear regime—as well as the influence of each CFG component—varies across classes, we empirically choose the interval for applying guidance and calculate the FDDINOv2 score with 50,000 generated images for each class separately. We summarize our observations as follows (see also section˜4.1):
@@
  • Role of CPC Guidance. CPC guidance generally improves generation quality, though its benefits can sometimes be less pronounced. For instance, in the tree frog and castle classes (Figure˜30), the CPC term does not enhance FDDINOv2 as much as the mean-shift term. Nevertheless, CPC guidance operates effectively over a wider range of guidance strengths $\gamma$ and noise intervals. For the green mamba and basketball classes, we show results within the prescribed noise interval as solid curves, and extend beyond this interval as dashed curves. While mean-shift becomes highly detrimental once outside the linear regime, CPC guidance remains beneficial.
@@ F.3 Mean-Shifted Noise Initialization
We perform the above experiments on 5 classes, where $\sigma(T)$ is set to 31.9. The results are shown in Figures˜32, 33, 34, 35 and 36. Note that the sample quality improves with a properly chosen $\gamma$.
@@ F.4 CFG in the Nonlinear Regime
We provide additional experimental results for section˜4.2 in Figure˜37. We argue that effective guidance in this regime should satisfy two key properties:
  • Capture local structure of a specific sample. As shown in Figure˜26, when $\sigma(t)$ is small, the model diverges considerably from its linear approximation, and linear CFG deviates from the actual nonlinear CFG. In this regime, CFG does not alter the overall image structure but instead refines existing features to produce crisper images. Consequently, effective guidance must adapt to each specific sample. We propose that such guidance can be derived from the network Jacobians $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t),\bm{c})$ evaluated at $\bm{x}_{t}$. Prior work [33] shows that the singular vectors of these Jacobians, which are equivalent to the posterior covariances, adapt to the input $\bm{x}_{t}$.
  • Capture class-specific patterns. As in the linear case, the guidance must also capture class-specific patterns. This can be achieved by contrasting the conditional Jacobian $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t),\bm{c})$ with the unconditional Jacobian $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t))$. Figure˜37 shows that guidance built using CPCs—i.e., the difference between these two Jacobians—yields effects similar to actual CFG. In contrast, guidance derived solely from the conditional Jacobian does not improve image quality.
@@
Appendix G Experimental Results on Latent Diffusion Models
In the main text, we conducted experiments using the EDM-1 model [4], which operates directly in pixel space with $64\times 64$ resolution. Here, we present complementary results on the EDM-2 [5] latent diffusion model, which generates images at $512\times 512$ resolution.
Linear Regime. We evaluate multiple guidance strategies—including actual CFG, linear CFG, Mean-shift guidance, positive CPC guidance, and negative CPC guidance—within the high-noise intervals (the linear regime). For each method, we generate 50,000 images conditioned on the class label “golden retriever” and compute the $\mathrm{FD_{DINOv2}}$ metric. The results, shown in Figure˜38, are consistent with the observations reported in the main text.
Nonlinear Regime. We next examine guidance effects in the nonlinear regime using (17) and (18). As shown in Figure˜39, guiding with CPCs produces visual effects similar to those of actual CFG—enhancing image sharpness and structure—whereas guidance with conditional PCs often leads to oversaturated colors. This highlights the importance of selectively amplifying class-specific features. We note that our heuristic guidance serves as a conceptual approximation and may not always perfectly align with actual CFG behavior. Additional failure cases will be provided in our code release.
Figure 38: Effects of CFG in the Linear Regime (EDM-2). Each row in (a) demonstrates the impact of different guidance types applied to EDM-2 within the linear regime (specified in the subtitles), with varying guidance strength $\gamma$. (b) shows the FDDINOv2 scores computed over 50,000 samples.
Figure 39: Effects of CFG in Nonlinear Regime (EDM-2). Different guidance methods, each with a fixed strength of $\gamma=8$, are applied at $\sigma(t)=1.502$. The samples in each row are generated from the same initial noise.
Appendix H CFG in Gaussian Mixture Model
@@
Appendix I Computing Resources